OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (zale101):

The equation of a linear function is: y = mx+b, m is the slope b is the y-intercept but there's another formula for finding the slope: Use the equation m = (y2-y1)/(x2-1) and get m = (2 - 1)/(- 4 - ( - 6)) = 1/2

OpenStudy (zale101):

can u solve from here?

OpenStudy (anonymous):

yes, thank you

OpenStudy (zale101):

np